# Enhancing Clinical Trial Patient Matching through Knowledge Augmentation and Reasoning with Multi-Agents

### v2

@@ 4.3. Experiment Settings
Our experiments were conducted on an Amazon EC2 g5.16xlarge instance, equipped with a A10G GPU with 24 GB of memory. Unless otherwise specified, all agents in the framework are powered by GPT-4o.
@@
For external knowledge augmentation, the Online Search Agent uses the Perplexity AI search engine to fetch up-to-date information from the web. The Retrieval Agent uses all-MiniLM-L12-v2, a pre-trained checkpoint from Sentence-Transformer (Reimers, 2019), as its encoder to construct the vector database of MedDRA and FDALabel.
Metric Zero-shot CoT MAKAR
Accuracy 0.8924 0.9201 0.9306
Precision 0.4615 0.5676 0.6333
Recall 0.6429 0.7500 0.6786
F1-Score 0.5373 0.6462 0.6552
Table 2. Trial-Level Comparison Across Different Methods
4.4. Results of Experiment 1
4.4.1. Criterion-Level Performance:
Table 1 shows the performance of three approaches—Zero-shot,CoT, MAKAR—across multiple individual criteria. Our results indicate that while the CoT method slightly outperforms Zero-shot in Accuracy and Recall, MAKAR consistently achieves higher scores, particularly in Accuracy (0.909) and Precision(0.822) on average. The F1-score of MAKAR (0.828) suggests a balanced performance across criteria, although it falls slightly behind in a few specific cases such as ”ABDOMINAL” and ”English”.
@@ 4.4.1. Criterion-Level Performance:
The MAKAR-augmented criteria, presented in Table LABEL:tab:aug_criteria, demonstrate significant improvements over the baseline human-redefined criteria outlined in Table LABEL:table:redefined_criteria from Wornow et al. (2024). These enhancements primarily focus on providing explicit clarifications and detailed expansions of key concepts, enabling LLMs to better align with trial requirements. For example, MAKAR supplements normal ranges for criteria like ”HBA1C” and ”CREATININE” across diverse demographic groups. And they are also the criteria that benefit the most from MAKAR’s augmentation. Additionally, MAKAR provides a more comprehensive definition for criteria like ”MAJOR-DIABETES,” and use more widely recognized terminology, ”heart attack”, to explain clinical terms ”Myocardial Infarction (MI)”. These refinements improve the ability of LLMs to match patients from clinical notes written in a variety of styles, resulting in notable performance gains for these criteria.
4.4.2. Trial-Level Performance
The result of trial-levle analysis can be found in Table 2. At the trial level, where a patient’s eligibility is determined by satisfying all selected criteria, MAKAR outperforms both Zero-shot and CoT. MAKAR achieves the highest Accuracy (0.9306), Precision (0.6333)and F1-score (0.6552) but underperform CoT on Recall. The trial-level results further validate MAKAR’s capability in maintaining a higher balance between identifying eligible patients and minimizing patient-filtering load
Inclusion Criteria
Metric GPT-4o o1 o3-mini CoT MAKAR
Accuracy 0.9000 0.9333 0.8533 0.9667 1.000
Precision 0.7884 0.8407 0.7211 0.9385 1.000
Recall 0.9020 0.9413 0.8151 0.9208 1.000
F1-Score 0.8281 0.8800 0.7516 0.9294 1.000
Exclusion Criteria
Metric GPT-4o o1 o3-mini CoT MAKARCorr
Accuracy 0.9000 0.9267 0.9133 0.9467 1.000
Precision 0.8693 0.8987 0.8840 0.9265 1.000
Recall 0.9090 0.9416 0.9253 0.9487 1.000
F1-Score 0.8845 0.9153 0.8999 0.9365 1.000
Table 3. Criterion-Level Comparison Across Different Methods in Experment 2
4.5. Results of Experiment 2
4.5.1. Results
Table 3 presents the evaluation results of our method on the in-house real-world dataset. We compared its performance against zero-shot prompting with GPT-4o, zero-shot prompting with two reasoning models (OpenAI-o3-mini and OpenAI-o1) and CoT strategy with GPT-4o. Our method achieves perfect scores across all four metrics.
Notably, for exclusion criteria, we report results for a corrected version of MAKAR (MAKARCorr). In the initial version, the model was not explicitly instructed on whether a given criterion was for inclusion or exclusion. Surprisingly, it was able to infer some exclusion criteria but focused on determining a patient’s overall eligibility for a trial rather than assessing whether a specific criterion was met. For instance, if a patient was 75 years old, they should have been marked as ”MET” for an exclusion criterion of ”greater than 70 years old”. While MAKAR correctly identified the age range and provided a logical reasoning process, it mistakenly output ”Not MET”—interpreting the task as determining overall trial eligibility rather than verifying the criterion itself. The corrected version addresses this by explicitly instructing the model to evaluate each criterion independently rather than making a broader eligibility judgment.
MAKAR outperforms all the benchmarks and the CoT follows closely, achieving 96.67% accuracy and 92.94% F1-score for inclusion criteria and 94.67% accuracy with 93.65% F1-score for exclusion criteria. o3-mini, a smaller model, is outperformed by o1 but shows competitive performance on exclusion criteria with zero-shot method with GPT-4o. Interestingly, Experiment 2 yields better overall performance than Experiment 1, despite its more knowledge-intensive criteria. We hypothesize that this is due to differences in the labeling process—the n2c2 challenge involves more subjective judgment, while our setup explicitly defines insufficient information as ”NOT MET”.
4.5.2. Error Analysis
Despite MAKAR achieving 100% accuracy in the second experiment, we manually checked the reasoning results of all 300 patient-criterion pair and identified two reasoning errors that did not affect the final output.
@@ 4.5.2. Error Analysis
First, in Exclusion Criterion 13 (Patients with ongoing indications for cardioprotective medication—ACE inhibitors, ARBs, and/or beta-blockers), MAKAR reasoned: ”The patient is on metoprolol, a beta-blocker, suggesting some ongoing indication for its use. However, without explicit information on the specific condition necessitating this medication (e.g., arrhythmia, prophylactic use post-cardiac event other than MI, or another reason), the justification remains uncertain.” This reasoning was overly conservative, but the Supervision Matching Agent correctly classified it as MET, aligning with expert judgment.
@@
Second, in Inclusion Criterion 5 (Histologically confirmed breast cancer with an invasive component $\geq$ 20 mm and/or morphologically confirmed spread to regional lymph nodes(stage cT2-cT4 with any cN, or cN1-cN3 with any cT).), MAKAR reasoned: ”The documentation mentions a suspicious breast nodule at the right breast (1.4 × 1.5 × 0.5 cm), which measures more than 20 mm in one dimension.” This was an obvious error, as the maximum dimension stated is only 1.5 cm. The mistake may come from over-reasoning, possibly computing the largest possible diagonal as $\sqrt{1.4^{2}+1.5^{2}+0.5^{2}}=2.1cm$. However, the final decision remained correct since the criterion lymph node involvement still holds.
4.6. Ablation
@@
In this section, we validate the contribution of each module in our framework. First, we remove the knowledge augmentation module to assess the reasoning module (MAKA-reason) using the original criteria, which essentially mirrors the CoT approach in Table 3. Next, we remove the reasoning module and evaluate the knowledge augmentation module (MAKA-aug) independently by performing zero-shot matching with the augmented criteria. The results, presented in Table 4, indicate that removing the reasoning module has a greater negative impact on inclusion criteria matching (F1-score: 88.37% vs. 92.94%), while removing the augmentation module has a larger effect on exclusion criteria (F1-score: 93.65% vs. 94.41%).
@@
Although we did not involve any reasoning-LLM-based agents, our approach still shows impressive reasoning capabilities. This can be attributed to the well-structured augmented criteria and the CoT strategy employed by the reasoning agent. For instance, when evaluating whether a patient meets the Inclusion Criterion 14 (Postmenopausal hormone receptor-positive patients), MAKA reasoned: ”…the record does not explicitly state whether the breast cancer was hormone receptor positive. However, the use of tamoxifen, which is specifically prescribed for estrogen receptor-positive breast cancer by blocking estrogen receptors, strongly implies that her breast cancer was hormone receptor positive.” Both MAKA and MAKA-reason successfully made the correct match, while MAKA-aug failed to capture this implicit relationship.
@@
Additionally, we also observed cases where knowledge augmentation improved performance. For example, in Exclusion Criterion 7, it mentions: ”…all histological lesions were HER2 1+ or 2+, not detected by FISH amplification, and HR positive according to ASCO guidelines”. MAKA-aug expanded the definitions from American Society of Clinical Oncology(ASCO) guidelines, made correct matches for all 10 patients. In contrast, MAKA-reason correctly identified only 8 out of 10.
Inclusion Criteria
Metric MAKAR-reason MAKAR-aug MAKAR
Accuracy 0.9667 0.9333 1.000
Precision 0.9385 0.8387 1.000
Recall 0.9208 0.9612 1.000
F1-Score 0.9294 0.8837 1.000
Exclusion Criteria
Metric MAKAR-reason MAKAR-aug MAKARCorr
Accuracy 0.9467 0.9533 1.000
Precision 0.9265 0.9360 1.000
Recall 0.9487 0.9534 1.000
F1-Score 0.9365 0.9441 1.000
Table 4. Results of Ablation Analysis
5. Discussion
@@
One interesting observation from our experiments is that the Router Agent consistently determined that all criteria could be augmented by the Self-Augmentation Agent, without requiring external retrieval or online search augmentation. This was true even for breast cancer-related criteria, which are typically complex and challenging for individuals without professional training.
@@
This relates to a limitation of our study, which comes from the dataset used for evaluation. Both the n2c2 dataset and the real-world dataset are relatively small and contain criteria that are not particularly challenging for LLM-based processing, allowing all benchmark methods to achieve strong performance. While MAKAR is designed as a general framework capable of handling both simple and complex criteria, these dataset constraints limit our ability to explore the knowledge boundaries of LLMs and fully assess the framework’s generalizability across a wider range of clinical trial requirements.
Moreover, this observation raises a discussion about the specific improvements our framework brings to baseline LLMs. While the framework is designed to bridge the knowledge gap between trial criteria and LLMs’ existing knowledge, our experiments indicate that LLMs can already expand and interpret the criteria in our datasets accurately. Instead, the primary benefit of knowledge augmentation lies in explicitly organizing relevant concepts and definitions in a well-structured format. By clearly outlining these key elements, the augmentation provides a type of breakdown of the subsequent reasoning steps, leading to improved overall performance.
@@
6. Conclusion
In this study, we introduced MAKAR, a scalable multi-agent framework designed to enhance patient-trial matching by combining knowledge augmentation with structured reasoning. By integrating an automated augmentation module with a reasoning module, MAKAR refines complex trial criteria and systematically evaluates patient eligibility. Our experimental results demonstrate that MAKAR outperforms existing benchmarks, including the state-of-the-art closed-source reasoning LLMs, and achieve 100% accuracy in a real-world offline test. And it also achieves up to an 8% improvement on complex criteria in a public dataset. Beyond performance gains, our findings emphasize the importance of explicitly structuring and expanding trial criteria to bridge the knowledge gap in AI-driven clinical reasoning.
It is worth noting that the design of MAKAR supports the future development of a human-AI collaboration system. For the trial-centric patient matching task, the workload of criterion-supervision performed by the critic agent is one-time per trial and relatively small and can be efficiently handled by human experts with minimal time investment. Exploring how humans can further enhance and facilitate this system presents an interesting direction for future research.
To address the limitations of this study and validate MAKAR’s generalizability, future work could also focus on evaluating the framework using more challenging datasets, such as those involving newly launched oncology trial that required up-to-date medical knowledge. Testing MAKAR in thess contexts would not only provide a more rigorous assessment of its capabilities but also demonstrate its potential to address critical real-world challenges in clinical trial recruitment.
7. Acknowledgments
We gratefully acknowledge support from the OpenAI Research Access Program, which provided access to advanced language models that were instrumental in developing and evaluating the framework presented in this work.
Appendix A Criteria
A.1. N2C2 Criteria Statistics
@@
Table 5. Counts of Eligible and Ineligible Patients for Each Criterion
A.2. N2C2 Original Criteria
Criteria Definition
DRUG-ABUSE Drug abuse, current or past.
ALCOHOL-ABUSE Current alcohol use over weekly recommended limits.
ENGLISH Patient must speak English.
MAKES-DECISIONS Patient must make their own medical decisions.
@@ A.2. N2C2 Original Criteria
ABDOMINAL History of intra-abdominal surgery, small or large intestine resection, or small bowel obstruction.
@@
MAJOR-DIABETES Major diabetes-related complications, such as: amputation, kidney damage, skin conditions, retinopathy, nephropathy, neuropathy.
@@
ADVANCED-CAD Advanced cardiovascular disease, defined as two or more of the following: medications for CAD, history of myocardial infarction, angina, ischemia.
MI-6MOS Myocardial infarction in the past 6 months.
KETO-1YR Diagnosis of ketoacidosis in the past year.
@@
DIETSUPP-2MOS Taken a dietary supplement (excluding Vitamin D) in the past 2 months.
ASP-FOR-MI Use of aspirin to prevent myocardial infarction.
HBA1C Any HbA1c value between 6.5 and 9.5%.
CREATININE Serum creatinine ¿ upper limit of normal.
Table 6. Definitions of the original criteria in the n2c2 dataset.
A.3. Redefined N2C2 Criteria
Table 7. Redefined Criteria proposed by (Wornow et al., 2024)
Criteria Definition
@@ A.3. Redefined N2C2 Criteria
ABDOMINAL History of intra-abdominal surgery. This could include any form of intra-abdominal surgery, including but not limited to small/large intestine resection or small bowel obstruction.
@@
ADVANCED-CAD Advanced cardiovascular disease (CAD). Defined as having 2 or more of the following: (a) Taking 2 or more medications to treat CAD (b) History of myocardial infarction (MI) (c) Currently experiencing angina (d) Ischemia, past or present. The patient must have at least 2 of these categories (a,b,c,d) to meet this criterion, otherwise the patient does not meet this criterion. For ADVANCED-CAD, be strict in your evaluation of the patient – if they just have cardiovascular disease, then they do not meet this criterion.
ALCOHOL-ABUSE Current alcohol use over weekly recommended limits.
ASP-FOR-MI Use of aspirin for preventing myocardial infarction (MI).
CREATININE Serum creatinine level above the upper normal limit.
@@
DIETSUPP-2MOS Consumption of a dietary supplement (excluding vitamin D) in the past 2 months. To assess this criterion, go through the list of medications and supplements taken from the note. If a substance could potentially be used as a dietary supplement (i.e., it is commonly used as a dietary supplement, even if it is not explicitly stated as being used as a dietary supplement), then the patient meets this criterion. Be lenient and broad in what is considered a dietary supplement. For example, a ’multivitamin’ and ’calcium carbonate’ should always be considered a dietary supplement if they are included in this list.
DRUG-ABUSE Current or past history of drug abuse.
@@
ENGLISH Patient speaks English. Assume that the patient speaks English, unless otherwise explicitly noted. If the patient’s language is not mentioned in the note, then assume they speak English and thus meet this criteria.
HBA1C Any hemoglobin A1c (HbA1c) value between 6.5% and 9.5%.
KETO-1YR Diagnosis of ketoacidosis within the past year.
@@
MAJOR-DIABETES Major diabetes-related complication. Examples of “major complication” (as opposed to “minor complication”) include, but are not limited to, any of the following that are a result of (or strongly correlated with) uncontrolled diabetes: • Amputation • Kidney damage • Skin conditions • Retinopathy • Nephropathy • Neuropathy. Additionally, if multiple conditions together imply a severe case of diabetes, then count that as a major complication.
@@
MAKES-DECISIONS Patient must make their own medical decisions. Assume that the patient makes their own medical decisions, unless otherwise explicitly noted. There is no information provided about the patient’s ability to make their own medical decisions, then assume they do make their own decisions and therefore meet this criteria.
MI-6MOS Myocardial infarction (MI) within the past 6 months.
A.4. MAKAR-Augemented N2C2 Criteria
Table 8. Redefined Criteria proposed by (Wornow et al., 2024)
Criteria Definition

### v3

@@ 4.3 Experiment Settings
Our experiments were conducted on 4 A10G GPU with 96 GB of memory. Unless otherwise specified, all agents in the framework are powered by GPT-4o.
@@
For external knowledge augmentation, the Online Search Agent uses the Perplexity AI search engine to fetch up-to-date information from the web. The Retrieval Agent uses all-MiniLM-L12-v2, a pre-trained checkpoint from Sentence-Transformer Reimers (2019), as its encoder to construct the vector database of MedDRA and FDALabel.
5 Results
Table 1 compares the performance of MAKAR with benchmark methods across two datasets. Overall, the MAKAR model consistently achieves the highest or near-highest accuracy and F1-scores across most conditions. MAKAR improves accuracy by approximately 4% and F1-score by around 7% on average. Notably, it shows particularly strong gains on criteria such as CREATININE, HBA1C, and MI-6MOS, achieving improvements of about 10% where other methods struggle.
@@ 5 Results
To better understand these gains, we examine the MAKAR-augmented criteria presented in Appendix B.3. Compared to the baseline human-refined criteria outlined in Appendix B.2, MAKAR has two major improvements: (1) more detailed expansions of key concepts, for example, supplementing normal ranges for criteria like HBA1C and CREATININE across diverse demographic groups; and (2) more comprehensive and accessible definitions, such as using widely recognized terms like “heart attack” alongside the clinical term “myocardial infarction (MI)”. These improvements improve the ability of LLMs to match patients from clinical notes written in various styles, resulting in substantial performance gains on these criteria.
Patient matching is not merely a binary classification task; the correctness of the reasoning process is essential for transparency, which in turn affects physicians’ trust in AI systems Asan et al. (2020). To evaluate this, we manually reviewed the reasoning outputs for all 300 patient–criterion pairs. Although MAKAR produced entirely correct final classifications on the ClinicalTrial dataset, we identified two minor flaws in the intermediate reasoning steps. And the results reported in Table 1 reflect the quality of both the reasoning process and the final matching outcomes.
@@
First, in Exclusion Criterion 13 (Patients with ongoing indications for cardioprotective medication—ACE inhibitors, ARBs, and/or beta-blockers), MAKAR reasoned: "The patient is on metoprolol, a beta-blocker, suggesting some ongoing indication for its use. However, without explicit information on the specific condition necessitating this medication (e.g., arrhythmia, prophylactic use post-cardiac event other than MI, or another reason), the justification remains uncertain.". Here, the Reasoning Agent is overly conservative, but the Supervision Matching Agent correctly classified this case as MET. This case shows the design is effective.
@@
Second,in Inclusion Criterion 5 (Histologically confirmed breast cancer with an invasive component $\geq$ 20 mm and/or morphologically confirmed spread to regional lymph nodes(stage cT2-cT4 with any cN, or cN1-cN3 with any cT).), MAKAR wrongly inferred that a 1.5 cm lesion is greater than 20 mm. However, because lymph node involvement was documented, the final decision remained correct.
Ablations Zero Shot MAKAR-Reason MAKAR-Aug MAKAR
F1-Score $0.857$ $0.928$ $0.933$ $0.982$
Reasoning Models
OpenAI o1 $0.898$
OpenAI o3 mini $0.825$
Table 2: Results of Ablation Analysis. We report only zero-shot results for reasoning models here, as they are not suited for CoT strategies Guo et al. (2025).
6 Ablation
6.1 Augmentation and Reasoning
@@
In this section, we validate the double-module design of our workflow on ClinicalTrial dataset. First, we remove the knowledge augmentation module to assess the reasoning module alone (MAKAR-Reason, equivalent to a CoT) using the original criteria described in Appendix B.1. Next, we remove the reasoning module and evaluate the knowledge augmentation module (MAKAR-Aug) using zero-shot matching with prompts detailed in Appendix C.1). The results, presented in Table 2, suggest that both modules contributes to overall performance. Specifically, removing the augmentation module (MAKAR-Reason alone) led to a 6% drop in F1 score, while removing the reasoning module (MAKAR-Aug alone) led to a 5% drop.
@@
To further demonstrate the effectiveness of the reasoning module, we compared its performance against other reasoning models. Interestingly, MAKAR-Reason outperformed both OpenAI o1 and o3 mini, which can be attributed to the chain-of-thought strategy employed by the reasoning agent. For example, when evaluating whether a patient met Inclusion Criterion 14 (Postmenopausal hormone receptor-positive patients), MAKAR reasoned: "…the record does not explicitly state whether the breast cancer was hormone receptor positive. However, the use of tamoxifen, which is specifically prescribed for estrogen receptor-positive breast cancer by blocking estrogen receptors, strongly implies that her breast cancer was hormone receptor positive." Both MAKAR and MAKAR-Reason correctly identified this case, while MAKAR-Aug failed to capture the implicit relationship.
@@
Additionally, we also observed cases where knowledge augmentation improved performance. For example, in Exclusion Criterion 7, it mentions: "…all histological lesions were HER2 1+ or 2+, not detected by FISH amplification, and HR positive according to ASCO guidelines". MAKA-aug expanded the definitions from American Society of Clinical Oncology(ASCO) guidelines, made correct matches for all 10 patients. In contrast, MAKA-reason correctly identified only 8 out of 10.
Underlying Models Zero Shot GPT-Aug MAKAR
Qwen 2.5 Instruct 7B $0.853$ $0.853$ $0.900$
Qwen 2.5 Instruct 14B $0.869$ $0.900$ $0.915$
LLaMA 3.1 Instruct 7B $0.815$ $0.815$ $0.853$
Gemma 3 Instruct 4B $0.788$ $0.788$ $0.828$
Gemma 3 Instruct 12B $0.860$ $0.869$ $0.884$
Mistral Instruct v0.3 7B $0.770$ $0.770$ $0.870$
GPT4o 0.900 0.943 0.987
Table 3: Results of Model Ablation in the Reasoning Module. GPT-Aug based on GPT-4o is equivalent to MAKAR-Aug.
6.2 Local Deployment
One key challenge for applying MAKAR in real-world settings is that agents in the reasoning module process patient EHR data. And such data is extremely sensitive and protected under The Health Insurance Portability and Accountability Act (HIPAA). It requires de-identifying large volumes of patient information before uploading data to a closed-source external model like GPT-4o. This process can be time-consuming and still have residual privacy risks. To address this limitation, we evaluated whether MAKAR can maintain high performance when replacing GPT-4o with open-source language models that can be deployed locally.
Our local deployment experiment consisted of three configurations: (1)Zero-Shot Matching, where we applied open-source models directly in zero-shot mode using the criteria described in Appendix B.4 and the prompts from Appendix C.1; (2) GPT-Aug, where we used GPT-4 to augment the criteria (since criteria text itself is not sensitive) and then applied the same zero-shot prompt with a open-source model for matching; and (3) MAKAR, where we replaced all agents in the reasoning module with open-source models to avoid uploading data to externally hosted servers.
The experiments were conducted on the ClinicalTrial dataset, and accuracies are reported in Table 3. We found that replacing GPT-4o in the reasoning module with locally deployed open-source models reduced performance by an average of 11%. However, some configurations still have competitive performance; for example, MAKAR powered by the Qwen family achieved accuracy greater than 0.900. These results further confirm the effectiveness of the MAKAR design, which improved performance by approximately 5% across open-source models.
Another interesting observation is that smaller models appeared less sensitive to augmented criteria: models under 7B parameters showed no performance gains from augmentation. In contrast, both smaller and larger models consistently benefited from the reasoning module, with accuracy improving after reasoning was applied. However, this observation should be interpreted with caution, as it is based on a limited number of model configurations evaluated in this study.
7 Related Work
This work falls within the literature on patient matching using free-text patient records. Early approaches such as RBC Oleynik et al. (2019) relied on rule-based classification. Transformer encoder-based methods were later introduced, including DeepEnroll Zhang et al. (2020) and COMPOSE Gao et al. (2020), which improved the ability to encode complex eligibility criteria and patient notes. More recently, generative large language models (GLLMs) have been applied to this task and have shown stronger performance, as in TrialGPT Jin et al. (2024), CoT matching Beattie et al. (2024), and other zero-shot methods Wornow et al. (2024). In parallel, Loaiza-Bonilla et al. (2025) explored the use of multi-agent systems to analyze oncology knowledge graphs for patient enrollment.
8 Conclusion
In this study, we introduced MAKAR, a scalable multi-agent workflow designed to enhance patient-trial matching by combining criterion augmentation with structured reasoning. Our experimental results show that MAKAR achieves an average improvement of 7% and up to 10% on certain criteria. Through detailed ablation studies, we further demonstrate that both modules of the MAKAR design contribute independently to these gains. In addition, our local deployment experiments confirmed that MAKAR remains effective when using open-source language models. Even with smaller models, such as Qwen and Gemma, the MAKAR workflow achieved competitive performance. These findings suggest that MAKAR can play a significant role in enabling more accurate and privacy-preserving patient selection.
Limitations
@@ 8 Conclusion
One observation from our experiments is that the Router Agent consistently determined that all criteria could be augmented by the Self-Augmentation Agent, without requiring external retrieval or online search augmentation. This was true even for breast cancer-related criteria, which are typically complex and challenging for individuals without professional training.
@@
This relates to a limitation of our study, which comes from the dataset used for evaluation. Both the n2c2 dataset and the real-world dataset are relatively small and contain criteria that are not particularly challenging for LLM-based processing, allowing all benchmark methods to achieve strong performance. While MAKAR is designed as a general framework capable of handling both simple and complex criteria, these dataset constraints limit our ability to explore the knowledge boundaries of LLMs and fully assess our method’s generalizability across a wider range of clinical trial requirements.
@@
Appendix A Statistics
@@
Table 4: Counts of Eligible and Ineligible Patients in N2C2
Appendix B Criteria
B.1 N2C2 Original Criteria
DRUG-ABUSE: Drug abuse, current or past.
ALCOHOL-ABUSE: Current alcohol use over weekly recommended limits.
ENGLISH: Patient must speak English.
MAKES-DECISIONS: Patient must make their own medical decisions.
@@ B.1 N2C2 Original Criteria
ABDOMINAL: History of intra-abdominal surgery, small or large intestine resection, or small bowel obstruction.
@@
MAJOR-DIABETES: Major diabetes-related complications, such as: amputation, kidney damage, skin conditions, retinopathy, nephropathy, neuropathy.
@@
ADVANCED-CAD: Advanced cardiovascular disease, defined as two or more of the following: medications for CAD, history of myocardial infarction, angina, ischemia.
MI-6MOS: Myocardial infarction in the past 6 months.
KETO-1YR: Diagnosis of ketoacidosis in the past year.
@@
DIETSUPP-2MOS: Taken a dietary supplement (excluding Vitamin D) in the past 2 months.
ASP-FOR-MI: Use of aspirin to prevent myocardial infarction.
HBA1C: Any HbA1c value between 6.5 and 9.5%.
CREATININE: Serum creatinine > upper limit of normal.
B.2 N2C2 Criteria Refined by Wornow et al. (2024)
@@
ABDOMINAL: History of intra-abdominal surgery. This could include any form of intra-abdominal surgery, including but not limited to small/large intestine resection or small bowel obstruction.
@@
ADVANCED-CAD: Advanced cardiovascular disease (CAD). For the purposes of this annotation, we define “advanced” as having 2 or more of the following: (a) Taking 2 or more medications to treat CAD (b) History of myocardial infarction (MI) (c) Currently experiencing angina (d) Ischemia, past or present. The patient must have at least 2 of these categories (a,b,c,d) to meet this criterion, otherwise the patient does not meet this criterion. For ADVANCED-CAD, be strict in your evaluation of the patient—if they just have cardiovascular disease, then they do not meet this criterion.
ALCOHOL-ABUSE: Current alcohol use over weekly recommended limits.
ASP-FOR-MI: Use of aspirin for preventing myocardial infarction (MI).
CREATININE: Serum creatinine level above the upper normal limit.
@@
DIETSUPP-2MOS: Consumption of a dietary supplement (excluding vitamin D) in the past 2 months. To assess this criterion, go through the list of medications and supplements taken from the note. If a substance could potentially be used as a dietary supplement (i.e., it is commonly used as a dietary supplement, even if it is not explicitly stated as being used as a dietary supplement), then the patient meets this criterion. Be lenient and broad in what is considered a dietary supplement. For example, a “multivitamin” and “calcium carbonate” should always be considered a dietary supplement if they are included in this list.
DRUG-ABUSE: Current or past history of drug abuse.
@@
ENGLISH: Patient speaks English. Assume that the patient speaks English, unless otherwise explicitly noted. If the patient’s language is not mentioned in the note, then assume they speak English and thus meet this criterion.
HBA1C: Any hemoglobin A1c (HbA1c) value between 6.5% and 9.5%.
KETO-1YR: Diagnosis of ketoacidosis within the past year.
@@
MAJOR-DIABETES: Major diabetes-related complication. Examples of “major complication” (as opposed to “minor complication”) include, but are not limited to, any of the following that are a result of (or strongly correlated with) uncontrolled diabetes: Amputation, kidney damage, skin conditions, retinopathy, nephropathy, neuropathy. Additionally, if multiple conditions together imply a severe case of diabetes, then count that as a major complication.
@@
MAKES-DECISIONS: Patient must make their own medical decisions. Assume that the patient makes their own medical decisions, unless otherwise explicitly noted. If there is no information provided about the patient’s ability to make their own medical decisions, then assume they do and therefore meet this criterion.
MI-6MOS: Myocardial infarction (MI) within the past 6 months.
B.3 N2C2 Criteria Refined by GPT-4o
ABDOMINAL